If X and Y are independent random variables with variances σ2X = 5 and σ2Y = 3, find the variance of the random variable Z = −2X +4Y − 3.
Repeat Exercise 4.62 if X and Y are not inde- pendent and σXY =1.
1
Expert's answer
2020-02-24T11:04:46-0500

If "X_1, X_2, ..., X_n" are independent random variables having normal distributions with means "\mu_1, \mu_2,..., \mu_n" and variances "\sigma_1^2, \sigma_2^2 , ..., \sigma_n ^2" respectively, then the random variable

"Y=a_1X_1+a_2X_2+...+a_nX_n" has a normal distribution with mean "\mu_Y=a_1\mu_1+a_2\mu_2+...+a_n\mu_n" and variance "\sigma_Y^2=a_1^2\sigma_1^2+a_2^2\sigma_2^2+...+a_n^2\sigma_n^2"

Then


"\sigma_Z^2=\sigma_{-2X+4Y-3}^2=(-2)^2\sigma_X^2+(4)^2\sigma_Y^2"

"\sigma_Z^2=\sigma_{-2X+4Y-3}^2=(-2)^2(5)+(4)^2(3)=68"

If "X" and "Y" are not independent then


"\sigma_{aX+bY+c}^2=a^2 \sigma_X^2+b^2\sigma_Y^2+2ab\sigma_{XY}"

"\sigma_Z^2=\sigma_{-2X+4Y-3}^2=(-2)^2(5)+(4)^2(3)+2(-2)(4)(1)=52"